# Card-Based Overwriting Protocol for Equality Function and Applications

### v1

@@ Abstract
Research in the area of secure multi-party computation with an unconventional method of using a physical deck of playing cards began in 1989 when den Boar proposed a protocol to compute the logical AND function using five cards. Since then, the area has gained interest from many researchers and several card-based protocols to compute various functions have been developed. In this paper, we propose a card-based protocol called the overwriting protocol that can securely compute the $k$-candidate $n$-variable equality function $f:\{0,1,\ldots,k-1\}^{n}\rightarrow\{0,1\}$. We also apply the technique used in this protocol to compute other similar functions.
@@ 1 Introduction
During a presidential election with $k$ candidates, a group of $n$ friends decide that they will talk about politics only if all of them support the same candidate. However, each person does not want to reveal his/her preference to the group (unless it is known that everyone supports the same candidate). They need a way to find out whether their preferences all coincide without leaking information of any individual’s preference (not even probabilistic information).
Theoretically, this is equivalent to each $i$-th person having an integer $a_{i}\in\{0,1,\ldots,k-1\}$, indicating the candidate he/she prefers. Define an equality function $E(a_{1},a_{2},\ldots,a_{n}):=1$ if $a_{1}=a_{2}=\ldots=a_{n}$, and $E(a_{1},a_{2},\ldots,a_{n}):=0$ otherwise. Our goal is to develop a protocol that can compute the value of $E(a_{1},a_{2},\ldots,a_{n})$ without leaking any other information.
@@ 1.1 Related Work
Research in card-based cryptography began in 1989 when den Boer [3] proposed a protocol called the five-card trick to compute the logical AND function of two players’ bits using five cards. After that, other AND function protocols [1, 2, 6, 7, 11, 12, 14, 19, 24] were also developed. These subsequently protocols either reduced the number of required cards or shuffles, or improved properties of the protocol involving output format, type of shuffles, etc.
@@ 1.1.1 Equality Function Protocols
Several previous results have been focused on computing the equality function. For the Boolean equality function (where $k=2$), a protocol for a special case $n=3$ has been independently proposed by Heather et al. [4] and by Shinagawa [22]. Later, Ruangwises and Itoh [20] developed two protocols to compute the Boolean equality function for any $n$. Their work can also be generalized to compute the $k$-candidate $n$-variable equality function $f:\{0,1,\ldots,k-1\}^{n}\rightarrow\{0,1\}$ for any $k$ and $n$ by converting each input into binary representation.
@@ 1.2 Our Contribution
In this paper, we propose a new card-based protocol called the overwriting protocol, which can securely compute the $k$-candidate $n$-variable equality function $f:\{0,1,\ldots,k-1\}^{n}\rightarrow\{0,1\}$ using $kn$ cards.
Comparing to the existing protocol of Ruangwises and Itoh [20] which uses $2\lceil\lg k\rceil n$ cards and $\lceil\lg k\rceil n-1$ shuffles, our protocol uses fewer cards when $k=3$ or 5 (and uses equal number of cards when $k=2$, 4, or 6), and also uses fewer shuffles for every $k\geq 3$. See Table 1 for comparison. In addition, our protocol is arguably simpler and more intuitive, as it does not require each player to convert an integer into binary representation.
@@
Set (§5.2) $S:\{0,1,\ldots,k-1\}^{n}\rightarrow P(\{0,1,\ldots,k-1\})$ $k(n+1)$ $n$
@@ 2.1 Equality Function
$E(a_{1},a_{2},\ldots,a_{n}):=\begin{cases}1,&\text{if $a_{1}=a_{2}=\ldots=a_{n}$;}\\ 0,&\text{otherwise.}\end{cases}$
@@ 3 Protocol for Equality Function
Figure 3: An example of a matrix $M$ constructed in Step 1, where $k=6$, $n=5$, $a_{1}=a_{3}=a_{5}=2$, $a_{2}=0$, and $a_{4}=3$ (with all cards actually being face-down)
@@ 5 Applications to Set and Set Size Functions
This situation is similar to that of the equality function where each $i$-th referee has an integer $a_{i}\in\{0,1,\ldots,k-1\}$ indicating the candidate he/she prefers, but we instead want to find a set of integers that are equal to at least one input value. Define a set function $S:\{0,1,\ldots,k-1\}^{n}\rightarrow P(\{0,1,\ldots,k-1\})$ as
$S(a_{1},a_{2},\ldots,a_{n}):=\{j|\exists i,a_{i}=j\}.$
For example, $S(3,2,3,0,5,0)=\{0,2,3,5\}$. Our goal is to develop a protocol to securely compute the set function between $n$ players.

### v2

@@ Abstract
Research in the area of secure multi-party computation with an unconventional method of using a physical deck of playing cards began in 1989 when den Boer proposed a protocol to compute the logical AND function using five cards. Since then, the area has gained interest from many researchers and several card-based protocols to compute various functions have been developed. In this paper, we propose a card-based protocol called the overwriting protocol that can securely compute the $k$-candidate $n$-variable equality function $f:\{0,1,\ldots,k-1\}^{n}\rightarrow\{0,1\}$. We also apply the technique used in this protocol to compute other similar functions.
@@ 1 Introduction
During a presidential election with $k$ candidates, a group of $n$ friends decide that they will talk about politics only if all of them support the same candidate. However, each person does not want to reveal his/her preference to the group (unless it is known that everyone supports the same candidate). They need a way to find out whether their preferences all coincide without leaking information about any individual’s preference (not even probabilistic information).
Theoretically, this is equivalent to each $i$-th person having an integer $a_{i}\in\{0,1,\ldots,k-1\}$, indicating the candidate he/she prefers. Define an equality function $E(a_{1},a_{2},\ldots,a_{n})\coloneqq 1$ if $a_{1}=a_{2}=\cdots=a_{n}$, and $E(a_{1},a_{2},\ldots,a_{n})\coloneqq 0$ otherwise. Our goal is to develop a protocol that can compute the value of $E(a_{1},a_{2},\ldots,a_{n})$ without leaking any other information.
@@ 1.1 Related Work
Research in card-based cryptography began in 1989 when den Boer [3] proposed a protocol called the five-card trick to compute the logical AND function of two players’ bits using five cards. After that, other AND function protocols [1, 2, 6, 7, 11, 12, 14, 19, 24] were also developed. These subsequent protocols either reduced the number of required cards or shuffles, or improved properties of the protocol involving output format, type of shuffles, etc.
@@ 1.1.1 Equality Function Protocols
Several previous results have been focused on computing the equality function. For the Boolean equality function (where $k=2$), a protocol for a special case $n=3$ has been independently proposed by Heather et al. [4] and by Shinagawa and Mizuki [22]. Later, Ruangwises and Itoh [20] developed two protocols to compute the Boolean equality function for any $n$. Their work can also be generalized to compute the $k$-candidate $n$-variable equality function $f:\{0,1,\ldots,k-1\}^{n}\rightarrow\{0,1\}$ for any $k$ and $n$ by converting each input into binary representation.
@@ 1.2 Our Contribution
In this paper, we propose a new card-based protocol called the overwriting protocol, which can securely compute the $k$-candidate $n$-variable equality function $f:\{0,1,\ldots,k-1\}^{n}\rightarrow\{0,1\}$ using $kn$ cards and $n$ shuffles.
Comparing to the existing protocol of Ruangwises and Itoh [20] which uses $2\lceil\lg k\rceil n$ cards and $\lceil\lg k\rceil n-1$ shuffles, our protocol uses fewer cards when $k=3$ or 5 (and uses equal number of cards when $k=2$, 4, or 6), and also uses fewer shuffles for every $k\geq 3$. See Table 1 for comparison. In addition, our protocol is simpler and more intuitive, as it does not require each player to convert an integer into binary representation.
@@
Set (§5.2) $S:\{0,1,\ldots,k-1\}^{n}\rightarrow\mathcal{P}(\{0,1,\ldots,k-1\})$ $k(n+1)$ $n$
@@ 2.1 Equality Function
$E(a_{1},a_{2},\ldots,a_{n})\coloneqq\begin{cases}1,&\text{if $a_{1}=a_{2}=\cdots=a_{n}$;}\\ 0,&\text{otherwise.}\end{cases}$
@@ 3 Protocol for Equality Function
Figure 3: An example of a matrix $M$ constructed in Step 1, where $k=6$, $n=5$, $a_{1}=a_{3}=a_{5}=2$, $a_{2}=3$, and $a_{4}=0$ (with all cards actually being face-down)
@@ 5 Applications to Set and Set Size Functions
This situation is similar to that of the equality function where each $i$-th referee has an integer $a_{i}\in\{0,1,\ldots,k-1\}$ indicating the candidate he/she prefers, but we instead want to find a set of integers that are equal to at least one input value. Define a set function $S:\{0,1,\ldots,k-1\}^{n}\rightarrow\mathcal{P}(\{0,1,\ldots,k-1\})$ as
$S(a_{1},a_{2},\ldots,a_{n})\coloneqq\{j|\exists i,a_{i}=j\},$
where $\mathcal{P}$ denotes the power set. For example, $S(3,2,3,0,5,0)=\{0,2,3,5\}$. Our goal is to develop a protocol to securely compute the set function between $n$ players.